# Are We Truly Forgetting? A Critical Re-examination of Machine Unlearning Evaluation Protocols

### v2

@@ Appendix C Additional Experiments
C.1 CKA Similarities between the Original Model and the Retrained Model.
We analyze the CKA similarities between $\theta_{o}$ and $\theta_{r}$ under different unlearning scenarios in Fig. 9. As expected, CKA similarities decrease under top class-wise forgetting scenarios and with increasing numbers of forgetting classes.
C.2 Backbone Architectures.
Table 3 provides an extended comparison of the baselines using alternative backbones, specifically ConvNeXt-T [25] and Swin-T [24]. We compare only model-agnostic methods.
Table 4: Comparison of logit-based evaluation across different unlearning baselines. Each block (Random, Office-Home Top, CUB Top, DomainNet-126 Top) indicates the subset of classes removed. For each subset, we report Accuracy on ImageNet-1K.
Random Office-Home Top CUB Top DomainNet-126 Top
# Unlearned classes Method FA RA TFA TRA Avg. Gap FA RA TFA TRA Avg. Gap FA RA TFA TRA Avg. Gap FA RA TFA TRA Avg. Gap
100 Retrained 0.0 (0.0) 76.0 (0.0) 0.0 (0.0) 75.6 (0.0) 0.0 0.0 (0.0) 76.2 (0.0) 0.0 (0.0) 76.1 (0.0) 0.0 0.0 (0.0) 74.9 (0.0) 0.0 (0.0) 74.3 (0.0) 0.0 0.0 (0.0) 76.4 (0.0) 0.0 (0.0) 75.9 (0.0) 0.0
FT 9.9 (9.9) 79.5 (3.5) 10.5 (10.5) 76.0 (0.4) 6.1 7.1 (7.1) 80.1 (4.1) 6.6 (6.6) 77.2 (1.1) 4.7 24.4 (24.4) 78.5 (3.6) 28.4 (28.4) 74.9 (0.6) 14.3 10.8 (10.8) 79.6 (3.2) 11.7 (11.7) 76.3 (0.4) 6.5
GA 1.5 (1.5) 11.4 (64.6) 1.5 (1.5) 11.3 (64.3) 33.0 0.5 (0.5) 0.2 (76.0) 0.4 (0.4) 0.2 (75.9) 38.2 0.1 (0.1) 3.6 (71.3) 0.0 (0.0) 3.8 (70.5) 35.5 0.0 (0.0) 0.4 (76.0) 0.0 (0.0) 0.5 (75.4) 37.7
RL 6.3 (6.3) 14.3 (61.7) 5.5 (5.5) 13.0 (62.6) 34.3 0.3 (0.3) 4.5 (71.7) 0.1 (0.1) 3.6 (72.5) 36.6 9.6 (9.6) 40.2 (34.7) 6.4 (6.4) 34.7 (39.6) 22.4 7.3 (7.3) 18.2 (58.2) 6.2 (6.2) 16.6 (59.3) 32.7
PL 1.0 (1.0) 79.5 (3.5) 1.0 (1.0) 76.9 (0.9) 1.6 1.1 (1.1) 80.0 (3.8) 1.0 (1.0) 77.9 (1.8) 2.0 0.7 (0.7) 78.6 (3.7) 0.6 (0.6) 75.5 (1.2) 1.6 1.1 (1.1) 79.5 (3.1) 1.0 (1.0) 77.0 (1.1) 1.6
SalUn 10.4 (10.4) 22.1 (53.9) 9.3 (9.3) 19.1 (56.5) 32.5 1.6 (1.6) 6.4 (69.8) 1.4 (1.4) 4.9 (71.2) 36.0 5.2 (5.2) 30.3 (44.6) 3.2 (3.2) 25.8 (48.5) 25.4 7.4 (7.4) 17.3 (59.1) 6.2 (6.2) 16.5 (59.4) 33.0
DUCK 0.9 (0.9) 74.6 (1.4) 0.9 (0.9) 74.5 (1.1) 1.1 1.5 (1.5) 73.9 (2.3) 1.4 (1.4) 74.4 (1.7) 1.7 6.8 (6.8) 74.7 (0.2) 8.5 (8.5) 73.6 (0.7) 4.1 1.7 (1.7) 76.8 (0.4) 2.1 (2.1) 76.0 (0.1) 1.1
CU 2.1 (2.1) 73.9 (2.1) 2.2 (2.2) 73.3 (2.3) 2.2 1.3 (1.3) 59.8 (16.4) 1.5 (1.5) 61.1 (15.0) 8.6 0.5 (0.5) 43.4 (31.5) 0.4 (0.4) 46.2 (28.1) 15.1 2.0 (2.0) 73.3 (3.1) 2.9 (2.9) 73.0 (2.9) 2.7
SCAR 4.2 (4.2) 80.5 (4.5) 3.9 (3.9) 77.4 (1.8) 3.6 5.0 (5.0) 81.3 (5.1) 5.0 (5.0) 78.7 (2.6) 4.4 29.0 (29.0) 76.1 (1.2) 33.4 (33.4) 74.6 (0.3) 16.0 13.0 (13.0) 77.1 (0.7) 13.4 (13.4) 76.0 (0.1) 6.8
SCRUB 1.1 (1.1) 67.3 (8.7) 1.1 (1.1) 65.7 (9.9) 5.2 0.0 (0.0) 0.1 (76.1) 0.0 (0.0) 0.1 (76.0) 38.0 0.0 (0.0) 3.5 (71.4) 0.0 (0.0) 4.2 (70.1) 35.4 0.0 (0.0) 2.0 (74.4) 0.0 (0.0) 2.8 (73.1) 36.9
200 Retrained 0.0 (0.0) 77.7 (0.0) 0.0 (0.0) 77.0 (0.0) 0.0 0.0 (0.0) 78.6 (0.0) 0.0 (0.0) 78.2 (0.0) 0.0 0.0 (0.0) 75.4 (0.0) 0.0 (0.0) 74.2 (0.0) 0.0 0.0 (0.0) 78.5 (0.0) 0.0 (0.0) 77.7 (0.0) 0.0
FT 11.7 (11.7) 80.7 (3.0) 13.0 (13.0) 77.2 (0.2) 7.0 21.1 (21.1) 84.4 (5.8) 20.1 (20.1) 67.1 (11.1) 14.5 44.5 (44.5) 82.0 (6.6) 46.9 (46.9) 77.0 (2.8) 25.2 17.3 (17.3) 84.1 (5.6) 17.9 (17.9) 80.4 (2.7) 10.9
GA 0.1 (0.1) 0.3 (77.4) 0.1 (0.1) 0.2 (76.8) 38.6 0.2 (0.2) 0.6 (78.0) 0.2 (0.2) 0.5 (77.7) 39.0 0.1 (0.1) 3.6 (71.8) 0.1 (0.1) 3.2 (71.0) 35.8 0.3 (0.3) 0.5 (78.0) 0.2 (0.2) 0.7 (77.0) 38.9
RL 4.4 (4.4) 8.0 (69.7) 3.8 (3.8) 6.4 (70.6) 37.1 15.7 (15.7) 33.9 (44.7) 14.6 (14.6) 31.9 (46.3) 30.3 17.1 (17.1) 42.1 (33.3) 13.2 (13.2) 36.2 (38.0) 25.4 19.1 (19.1) 29.8 (48.7) 17.9 (17.9) 27.8 (49.9) 33.9
PL 1.2 (1.2) 80.2 (2.5) 1.0 (1.0) 78.1 (1.1) 1.5 1.1 (1.1) 81.1 (2.5) 1.0 (1.0) 79.5 (1.3) 1.5 0.9 (0.9) 78.2 (2.8) 0.9 (0.9) 75.4 (1.2) 1.5 1.2 (1.2) 80.6 (2.1) 1.0 (1.0) 78.9 (1.2) 1.4
SalUn 9.7 (9.7) 14.4 (63.3) 8.5 (8.5) 12.6 (64.4) 36.5 7.9 (7.9) 18.3 (60.3) 7.3 (7.3) 16.8 (61.4) 34.2 3.3 (3.3) 14.7 (60.7) 2.5 (2.5) 12.2 (62.0) 32.1 9.5 (9.5) 14.0 (64.5) 9.1 (9.1) 12.8 (64.9) 37.0
DUCK 0.8 (0.8) 78.1 (0.4) 0.8 (0.8) 77.3 (0.3) 0.6 1.0 (1.0) 78.4 (0.2) 1.2 (1.2) 78.4 (0.2) 0.7 5.7 (5.7) 74.8 (0.6) 7.0 (7.0) 73.7 (0.5) 3.5 0.7 (0.7) 78.4 (0.1) 1.0 (1.0) 77.8 (0.1) 0.5
CU 1.0 (1.0) 79.7 (2.0) 1.2 (1.2) 77.4 (0.4) 1.2 0.8 (0.8) 79.5 (0.9) 1.1 (1.1) 78.0 (0.2) 0.8 3.4 (3.4) 75.5 (0.1) 4.6 (4.6) 73.5 (0.7) 2.2 1.4 (1.4) 75.7 (2.8) 1.9 (1.9) 75.4 (2.3) 2.1
SCAR 15.0 (15.0) 80.8 (3.1) 15.6 (15.6) 78.7 (1.7) 8.9 13.7 (13.7) 79.1 (0.5) 13.7 (13.7) 78.4 (0.2) 7.0 35.5 (35.5) 75.9 (0.5) 37.1 (37.1) 74.4 (0.2) 18.3 11.3 (11.3) 78.8 (0.3) 12.5 (12.5) 78.1 (0.4) 6.1
SCRUB 18.3 (18.3) 72.5 (5.2) 16.9 (16.9) 69.7 (7.3) 11.9 0.0 (0.0) 0.1 (78.5) 0.0 (0.0) 0.1 (78.1) 39.2 0.0 (0.0) 3.9 (71.5) 0.0 (0.0) 5.0 (69.2) 35.2 0.2 (0.2) 1.2 (77.3) 0.3 (0.3) 1.8 (75.9) 38.4
300 Retrained 0.0 (0.0) 78.9 (0.0) 0.0 (0.0) 77.6 (0.0) 0.0 0.0 (0.0) 80.7 (0.0) 0.0 (0.0) 79.9 (0.0) 0.0 0.0 (0.0) 76.9 (0.0) 0.0 (0.0) 75.1 (0.0) 0.0 0.0 (0.0) 80.3 (0.0) 0.0 (0.0) 79.7 (0.0) 0.0
FT 14.0 (14.0) 82.0 (3.1) 15.4 (15.4) 78.9 (1.3) 8.5 28.4 (28.4) 85.3 (4.6) 28.6 (28.6) 82.3 (2.4) 16.0 45.5 (45.5) 82.7 (5.8) 46.7 (46.7) 77.7 (2.6) 25.2 21.7 (21.7) 85.4 (5.1) 22.3 (22.3) 82.3 (2.6) 12.9
GA 0.3 (0.3) 0.3 (78.6) 0.2 (0.2) 0.3 (77.3) 39.1 0.2 (0.2) 0.2 (80.5) 0.2 (0.2) 0.3 (79.6) 40.1 0.2 (0.2) 2.0 (74.9) 0.2 (0.2) 1.8 (73.3) 37.2 0.2 (0.2) 0.3 (80.0) 0.2 (0.2) 0.4 (79.3) 39.9
RL 6.4 (6.4) 9.0 (69.9) 5.5 (5.5) 8.1 (69.5) 37.8 11.5 (11.5) 22.9 (57.8) 10.2 (10.2) 21.2 (58.7) 34.6 11.9 (11.9) 24.1 (52.7) 9.2 (9.2) 20.7 (54.4) 32.0 14.4 (14.4) 19.4 (60.9) 13.3 (13.3) 18.0 (61.7) 37.6
PL 1.0 (1.0) 81.1 (2.2) 0.9 (0.9) 79.3 (1.7) 1.5 1.0 (1.0) 81.9 (1.2) 1.0 (1.0) 81.0 (1.0) 1.1 1.0 (1.0) 78.6 (1.7) 0.9 (0.9) 76.0 (0.9) 1.1 1.0 (1.0) 81.9 (1.6) 0.8 (0.8) 80.9 (1.2) 1.2
SalUn 10.5 (10.5) 14.3 (64.6) 8.6 (8.6) 12.4 (65.2) 37.2 8.3 (8.3) 14.7 (66.0) 8.2 (8.2) 13.4 (66.5) 37.3 7.3 (7.3) 17.0 (59.9) 6.0 (6.0) 15.3 (59.8) 33.3 10.5 (10.5) 14.5 (65.8) 9.5 (9.5) 13.4 (66.3) 38.0
DUCK 0.9 (0.9) 79.5 (0.6) 1.0 (1.0) 75.6 (2.0) 1.1 2.4 (2.4) 80.2 (0.5) 2.7 (2.7) 80.0 (0.1) 1.4 9.3 (9.3) 76.5 (0.4) 11.5 (11.5) 74.6 (0.5) 5.4 1.7 (1.7) 80.4 (0.1) 1.9 (1.9) 80.3 (0.6) 1.1
CU 0.7 (0.7) 81.1 (2.2) 0.8 (0.8) 78.6 (1.0) 1.2 1.6 (1.6) 81.5 (0.8) 2.0 (2.0) 79.9 (0.0) 1.1 6.7 (6.7) 77.9 (1.0) 8.8 (8.8) 74.9 (0.2) 4.2 1.0 (1.0) 82.0 (1.7) 1.2 (1.2) 80.7 (1.0) 1.2
SCAR 13.7 (13.7) 81.7 (2.8) 14.6 (14.6) 80.0 (2.4) 8.4 20.8 (20.8) 82.7 (2.0) 21.0 (21.0) 81.7 (1.8) 11.4 40.5 (40.5) 79.6 (2.7) 41.1 (41.1) 76.9 (1.8) 21.4 15.1 (15.1) 82.7 (2.4) 15.6 (15.6) 81.8 (2.1) 8.8
SCRUB 55.6 (55.6) 80.9 (2.0) 51.7 (51.7) 77.5 (0.1) 27.4 0.1 (0.1) 0.1 (80.6) 0.1 (0.1) 0.1 (79.8) 40.2 0.2 (0.2) 3.9 (73.0) 0.2 (0.2) 4.5 (70.6) 36.0 1.1 (1.1) 12.9 (67.4) 1.1 (1.1) 15.6 (64.1) 33.4
Table 5: Comparison of $k$-NN performance across different unlearning subsets. Each block (Random, Office-Home Top, CUB Top, DomainNet-126 Top) indicates the subset of classes removed. For each subset, we report $k$-NN Accuracy on three downstream tasks (Office-Home, CUB, and DomainNet-126) to gauge transferability.
Random Office-Home Top CUB Top DomainNet-126 Top
# Unlearned classes Method Office-Home $k$-NN CUB $k$-NN DomainNet-126 $k$-NN Office-Home $k$-NN CUB $k$-NN DomainNet-126 $k$-NN
@@
Table 6: Comparison of CKA values bewteen $\theta_{u}$ and $\theta_{r}$ across different unlearning subsets. Each block (Random, Office-Home Top, CUB Top, DomainNet-126 Top) indicates the subset of classes removed. For each subset, we report CKA value on three downstream tasks (Office-Home, CUB, and DomainNet-126) to gauge representation quality.
Random Office-Home Top CUB Top DomainNet-126 Top
Unlearned classes Method Office-Home CKA CUB CKA DomainNet-126 CKA Avg. CKA Office-Home CKA CUB CKA DomainNet-126 CKA
@@
Table 7: Comparison of the proposed metrics across different unlearning subsets. Each block (Random, Office-Home Top, CUB Top, DomainNet-126 Top) indicates the subset of classes removed. The evaluated metrics include AGL, AGR, and the unified metric H-LR. Higher AGL, AGR, and H-LR scores indicate better preservation of feature representations and overall unlearning effectiveness.
Random Office-Home Top CUB Top DomainNet-126 Top
# Unlearned classes Method AGL AGR H-LR AGL AGR H-LR AGL AGR H-LR AGL AGR H-LR

### v3

@@ Appendix B Additional Experimental Results
Table 7: Unlearning evaluation with 100 Random Class-wise Forgetting on ConvNeXt and Swin Transformers. The bold numbers represent optimal values.
Standard deviations. The standard deviations are in Table 8 to quantify the variability and uncertainty of the measured results.
ImageNet-1K Office-Home CUB DomainNet Office-Home CUB DomainNet
Method FA RA TFA TRA $k$-NN CKA($\theta_{u}$, $\theta_{r}$)
[…]
Table 8: The standard deviation of each value in Tables 2, 3 in the main paper. The standard deviations correspond to 95% confidence intervals, with each algorithm run 5 times.
t-SNE visualizations. Fig. 12 provides a comprehensive t-SNE visualizations for all baseline methods, expanding upon the subset shown in Figure 3 in the main text. These plots visualize the feature embeddings from the encoder for a consistent subset of retain and forget classes (red).
(a) Original ($\theta_{o}$)
(b) Retrained ($\theta_{r}$)
(c) FT
(d) GA
(e) RL
(f) SalUn
(g) PL
(h) DUCK
(i) CU
(j) SCAR
(k) SCRUB
Impact of scaling the number of forgetting classes. The full results in Tables 915 allow for a detailed analysis of algorithm scalability as the number of forgetting classes (N) increases from 100 to 300. This analysis reveals a critical divergence in stability. As detailed by the H-LR scores in Table 15, methods like PL, DUCK, CU demonstrate remarkable robustness. For instance, in the Random scenario, PL’s H-LR score only drops slightly from 0.89 (N=100) to 0.86 (N=300). In stark contrast, several methods that rely on more complex optimization objectives suffer a performance collapse. SCRUB’s H-LR score plummets from 0.69 (N=100) to a mere 0.33 (N=300). This scalability failure reinforces our hypothesis from Section 4.4 about “unbounded loss functions” and optimization instability. It provides empirical evidence that these methods are not practically scalable for large-scale unlearning tasks, despite their promising results in small-scale or small-N settings.
Figure 13: Comparison of PL baseline and PL with Nesterov Momentum.
Impact of optimization strategies on shortcut learning. In Section 3.7 (“Why does the unlearned model maintain high similarity to the original model?”), we discussed the phenomenon of “shortcut learning”, where the unlearning process merely modifies the classifier. Addressing the hypothesis that optimization dynamics might be the cause, we investigated whether acceleration methods, specifically Nesterov momentum, could enable the model to bypass these sub-optimal solutions [24]. We compared the Pseudo Labeling (PL) baseline against a version optimized with Nesterov momentum under the Random-100 Class-wise Forgetting scenario. As illustrated in Fig. 13, the inclusion of momentum resulted in negligible differences across both logit-based (FA, RA) and representation-based ($CKA(\theta_{u},\theta_{r})$) metrics. These results indicate that conventional momentum strategies alone cannot mitigate the shortcut learning embedded in current unlearning objectives. This suggests that addressing shortcut learning may require broader changes—such as revisiting the objective formulation or modifying architectural components. We leave the exploration of more advanced strategies to future work.
Random
Classes Method FA RA TFA TRA Avg. Gap
100 Retrained 0.0 (0.0) 76.0 (0.0) 0.0 (0.0) 75.6 (0.0) 0.0
FT 9.9 (9.9) 79.5 (3.5) 10.5 (10.5) 76.0 (0.4) 6.1
GA 1.5 (1.5) 11.4 (64.6) 1.5 (1.5) 11.3 (64.3) 33.0
RL 6.3 (6.3) 14.3 (61.7) 5.5 (5.5) 13.0 (62.6) 34.3
PL 1.0 (1.0) 79.5 (3.5) 1.0 (1.0) 76.9 (0.9) 1.6
SalUn 10.4 (10.4) 22.1 (53.9) 9.3 (9.3) 19.1 (56.5) 32.5
DUCK 0.9 (0.9) 74.6 (1.4) 0.9 (0.9) 74.5 (1.1) 1.1
CU 2.1 (2.1) 73.9 (2.1) 2.2 (2.2) 73.3 (2.3) 2.2
SCAR 4.2 (4.2) 80.5 (4.5) 3.9 (3.9) 77.4 (1.8) 3.6
SCRUB 1.1 (1.1) 67.3 (8.7) 1.1 (1.1) 65.7 (9.9) 5.2
200 Retrained 0.0 (0.0) 77.7 (0.0) 0.0 (0.0) 77.0 (0.0) 0.0
FT 11.7 (11.7) 80.7 (3.0) 13.0 (13.0) 77.2 (0.2) 7.0
GA 0.1 (0.1) 0.3 (77.4) 0.1 (0.1) 0.2 (76.8) 38.6
RL 4.4 (4.4) 8.0 (69.7) 3.8 (3.8) 6.4 (70.6) 37.1
PL 1.2 (1.2) 80.2 (2.5) 1.0 (1.0) 78.1 (1.1) 1.5
SalUn 9.7 (9.7) 14.4 (63.3) 8.5 (8.5) 12.6 (64.4) 36.5
DUCK 0.8 (0.8) 78.1 (0.4) 0.8 (0.8) 77.3 (0.3) 0.6
CU 1.0 (1.0) 79.7 (2.0) 1.2 (1.2) 77.4 (0.4) 1.2
SCAR 15.0 (15.0) 80.8 (3.1) 15.6 (15.6) 78.7 (1.7) 8.9
SCRUB 18.3 (18.3) 72.5 (5.2) 16.9 (16.9) 69.7 (7.3) 11.9
300 Retrained 0.0 (0.0) 78.9 (0.0) 0.0 (0.0) 77.6 (0.0) 0.0
FT 14.0 (14.0) 82.0 (3.1) 15.4 (15.4) 78.9 (1.3) 8.5
GA 0.3 (0.3) 0.3 (78.6) 0.2 (0.2) 0.3 (77.3) 39.1
RL 6.4 (6.4) 9.0 (69.9) 5.5 (5.5) 8.1 (69.5) 37.8
PL 1.0 (1.0) 81.1 (2.2) 0.9 (0.9) 79.3 (1.7) 1.5
SalUn 10.5 (10.5) 14.3 (64.6) 8.6 (8.6) 12.4 (65.2) 37.2
DUCK 0.9 (0.9) 79.5 (0.6) 1.0 (1.0) 75.6 (2.0) 1.1
CU 0.7 (0.7) 81.1 (2.2) 0.8 (0.8) 78.6 (1.0) 1.2
SCAR 13.7 (13.7) 81.7 (2.8) 14.6 (14.6) 80.0 (2.4) 8.4
SCRUB 55.6 (55.6) 80.9 (2.0) 51.7 (51.7) 77.5 (0.1) 27.4
Table 9: Random class-wise Forgetting results. The blue numbers indicate the gap. The bold numbers represent optimal values.
Office-Home Top
Classes Method FA RA TFA TRA Avg. Gap
100 Retrained 0.0 (0.0) 76.2 (0.0) 0.0 (0.0) 76.1 (0.0) 0.0
FT 7.1 (7.1) 80.1 (4.1) 6.6 (6.6) 77.2 (1.1) 4.7
GA 0.5 (0.5) 0.2 (76.0) 0.4 (0.4) 0.2 (75.9) 38.2
RL 0.3 (0.3) 4.5 (71.7) 0.1 (0.1) 3.6 (72.5) 36.6
PL 1.1 (1.1) 80.0 (3.8) 1.0 (1.0) 77.9 (1.8) 2.0
SalUn 1.6 (1.6) 6.4 (69.8) 1.4 (1.4) 4.9 (71.2) 36.0
DUCK 1.5 (1.5) 73.9 (2.3) 1.4 (1.4) 74.4 (1.7) 1.7
CU 1.3 (1.3) 59.8 (16.4) 1.5 (1.5) 61.1 (15.0) 8.6
SCAR 5.0 (5.0) 81.3 (5.1) 5.0 (5.0) 78.7 (2.6) 4.4
SCRUB 0.0 (0.0) 0.1 (76.1) 0.0 (0.0) 0.1 (76.0) 38.0
200 Retrained 0.0 (0.0) 78.6 (0.0) 0.0 (0.0) 78.2 (0.0) 0.0
FT 21.1 (21.1) 84.4 (5.8) 20.1 (20.1) 67.1 (11.1) 14.5
GA 0.2 (0.2) 0.6 (78.0) 0.2 (0.2) 0.5 (77.7) 39.0
RL 15.7 (15.7) 33.9 (44.7) 14.6 (14.6) 31.9 (46.3) 30.3
PL 1.1 (1.1) 81.1 (2.5) 1.0 (1.0) 79.5 (1.3) 1.5
SalUn 7.9 (7.9) 18.3 (60.3) 7.3 (7.3) 16.8 (61.4) 34.2
DUCK 1.0 (1.0) 78.4 (0.2) 1.2 (1.2) 78.4 (0.2) 0.7
CU 0.8 (0.8) 79.5 (0.9) 1.1 (1.1) 78.0 (0.2) 0.8
SCAR 13.7 (13.7) 79.1 (0.5) 13.7 (13.7) 78.4 (0.2) 7.0
SCRUB 0.0 (0.0) 0.1 (78.5) 0.0 (0.0) 0.1 (78.1) 39.2
300 Retrained 0.0 (0.0) 80.7 (0.0) 0.0 (0.0) 79.9 (0.0) 0.0
FT 28.4 (28.4) 85.3 (4.6) 28.6 (28.6) 82.3 (2.4) 16.0
GA 0.2 (0.2) 0.2 (80.5) 0.2 (0.2) 0.3 (79.6) 40.1
RL 11.5 (11.5) 22.9 (57.8) 10.2 (10.2) 21.2 (58.7) 34.6
PL 1.0 (1.0) 81.9 (1.2) 1.0 (1.0) 81.0 (1.0) 1.1
SalUn 8.3 (8.3) 14.7 (66.0) 8.2 (8.2) 13.4 (66.5) 37.3
DUCK 2.4 (2.4) 80.2 (0.5) 2.7 (2.7) 80.0 (0.1) 1.4
CU 1.6 (1.6) 81.5 (0.8) 2.0 (2.0) 79.9 (0.0) 1.1
SCAR 20.8 (20.8) 82.7 (2.0) 21.0 (21.0) 81.7 (1.8) 11.4
SCRUB 0.1 (0.1) 0.1 (80.6) 0.1 (0.1) 0.1 (79.8) 40.2
Table 10: Office-Home Top Class-wise Forgetting results. The blue numbers indicate the gap. The bold numbers represent optimal values.
CUB Top
Classes Method FA RA TFA TRA Avg. Gap
100 Retrained 0.0 (0.0) 74.9 (0.0) 0.0 (0.0) 74.3 (0.0) 0.0
FT 24.4 (24.4) 78.5 (3.6) 28.4 (28.4) 74.9 (0.6) 14.3
GA 0.1 (0.1) 3.6 (71.3) 0.0 (0.0) 3.8 (70.5) 35.5
RL 9.6 (9.6) 40.2 (34.7) 6.4 (6.4) 34.7 (39.6) 22.4
PL 0.7 (0.7) 78.6 (3.7) 0.6 (0.6) 75.5 (1.2) 1.6
SalUn 5.2 (5.2) 30.3 (44.6) 3.2 (3.2) 25.8 (48.5) 25.4
DUCK 6.8 (6.8) 74.7 (0.2) 8.5 (8.5) 73.6 (0.7) 4.1
CU 0.5 (0.5) 43.4 (31.5) 0.4 (0.4) 46.2 (28.1) 15.1
SCAR 29.0 (29.0) 76.1 (1.2) 33.4 (33.4) 74.6 (0.3) 16.0
SCRUB 0.0 (0.0) 3.5 (71.4) 0.0 (0.0) 4.2 (70.1) 35.4
200 Retrained 0.0 (0.0) 75.4 (0.0) 0.0 (0.0) 74.2 (0.0) 0.0
FT 44.5 (44.5) 82.0 (6.6) 46.9 (46.9) 77.0 (2.8) 25.2
GA 0.1 (0.1) 3.6 (71.8) 0.1 (0.1) 3.2 (71.0) 35.8
RL 17.1 (17.1) 42.1 (33.3) 13.2 (13.2) 36.2 (38.0) 25.4
PL 0.9 (0.9) 78.2 (2.8) 0.9 (0.9) 75.4 (1.2) 1.5
SalUn 3.3 (3.3) 14.7 (60.7) 2.5 (2.5) 12.2 (62.0) 32.1
DUCK 5.7 (5.7) 74.8 (0.6) 7.0 (7.0) 73.7 (0.5) 3.5
CU 3.4 (3.4) 75.5 (0.1) 4.6 (4.6) 73.5 (0.7) 2.2
SCAR 35.5 (35.5) 75.9 (0.5) 37.1 (37.1) 74.4 (0.2) 18.3
SCRUB 0.0 (0.0) 3.9 (71.5) 0.0 (0.0) 5.0 (69.2) 35.2
300 Retrained 0.0 (0.0) 76.9 (0.0) 0.0 (0.0) 75.1 (0.0) 0.0
FT 45.5 (45.5) 82.7 (5.8) 46.7 (46.7) 77.7 (2.6) 25.2
GA 0.2 (0.2) 2.0 (74.9) 0.2 (0.2) 1.8 (73.3) 37.2
RL 11.9 (11.9) 24.1 (52.7) 9.2 (9.2) 20.7 (54.4) 32.0
PL 1.0 (1.0) 78.6 (1.7) 0.9 (0.9) 76.0 (0.9) 1.1
SalUn 7.3 (7.3) 17.0 (59.9) 6.0 (6.0) 15.3 (59.8) 33.3
DUCK 9.3 (9.3) 76.5 (0.4) 11.5 (11.5) 74.6 (0.5) 5.4
CU 6.7 (6.7) 77.9 (1.0) 8.8 (8.8) 74.9 (0.2) 4.2
SCAR 40.5 (40.5) 79.6 (2.7) 41.1 (41.1) 76.9 (1.8) 21.4
SCRUB 0.2 (0.2) 3.9 (73.0) 0.2 (0.2) 4.5 (70.6) 36.0
Table 11: CUB Top Class-wise Forgetting results. The blue numbers indicate the gap. The bold numbers represent optimal values.
DomainNet Top
Classes Method FA RA TFA TRA Avg. Gap
100 Retrained 0.0 (0.0) 76.4 (0.0) 0.0 (0.0) 75.9 (0.0) 0.0
FT 10.8 (10.8) 79.6 (3.2) 11.7 (11.7) 76.3 (0.4) 6.5
GA 0.0 (0.0) 0.4 (76.0) 0.0 (0.0) 0.5 (75.4) 37.7
RL 7.3 (7.3) 18.2 (58.2) 6.2 (6.2) 16.6 (59.3) 32.7
PL 1.1 (1.1) 79.5 (3.1) 1.0 (1.0) 77.0 (1.1) 1.6
SalUn 7.4 (7.4) 17.3 (59.1) 6.2 (6.2) 16.5 (59.4) 33.0
DUCK 1.7 (1.7) 76.8 (0.4) 2.1 (2.1) 76.0 (0.1) 1.1
CU 2.0 (2.0) 73.3 (3.1) 2.9 (2.9) 73.0 (2.9) 2.7
SCAR 13.0 (13.0) 77.1 (0.7) 13.4 (13.4) 76.0 (0.1) 6.8
SCRUB 0.0 (0.0) 2.0 (74.4) 0.0 (0.0) 2.8 (73.1) 36.9
200 Retrained 0.0 (0.0) 78.5 (0.0) 0.0 (0.0) 77.7 (0.0) 0.0
FT 17.3 (17.3) 84.1 (5.6) 17.9 (17.9) 80.4 (2.7) 10.9
GA 0.3 (0.3) 0.5 (78.0) 0.2 (0.2) 0.7 (77.0) 38.9
RL 19.1 (19.1) 29.8 (48.7) 17.9 (17.9) 27.8 (49.9) 33.9
PL 1.2 (1.2) 80.6 (2.1) 1.0 (1.0) 78.9 (1.2) 1.4
SalUn 9.5 (9.5) 14.0 (64.5) 9.1 (9.1) 12.8 (64.9) 37.0
DUCK 0.7 (0.7) 78.4 (0.1) 1.0 (1.0) 77.8 (0.1) 0.5
CU 1.4 (1.4) 75.7 (2.8) 1.9 (1.9) 75.4 (2.3) 2.1
SCAR 11.3 (11.3) 78.8 (0.3) 12.5 (12.5) 78.1 (0.4) 6.1
SCRUB 0.2 (0.2) 1.2 (77.3) 0.3 (0.3) 1.8 (75.9) 38.4
300 Retrained 0.0 (0.0) 80.3 (0.0) 0.0 (0.0) 79.7 (0.0) 0.0
FT 21.7 (21.7) 85.4 (5.1) 22.3 (22.3) 82.3 (2.6) 12.9
GA 0.2 (0.2) 0.3 (80.0) 0.2 (0.2) 0.4 (79.3) 39.9
RL 14.4 (14.4) 19.4 (60.9) 13.3 (13.3) 18.0 (61.7) 37.6
PL 1.0 (1.0) 81.9 (1.6) 0.8 (0.8) 80.9 (1.2) 1.2
SalUn 10.5 (10.5) 14.5 (65.8) 9.5 (9.5) 13.4 (66.3) 38.0
DUCK 1.7 (1.7) 80.4 (0.1) 1.9 (1.9) 80.3 (0.6) 1.1
CU 1.0 (1.0) 82.0 (1.7) 1.2 (1.2) 80.7 (1.0) 1.2
SCAR 15.1 (15.1) 82.7 (2.4) 15.6 (15.6) 81.8 (2.1) 8.8
SCRUB 1.1 (1.1) 12.9 (67.4) 1.1 (1.1) 15.6 (64.1) 33.4
Table 12: DomainNet Top Class-wise Forgetting results. The blue numbers indicate the gap. The bold numbers represent optimal values.
Random Office-Home Top CUB Top DomainNet Top
Classes Method Office-Home $k$-NN CUB $k$-NN DomainNet $k$-NN Office-Home $k$-NN CUB $k$-NN DomainNet $k$-NN
@@
Table 13: Comparison of $k$-NN performance across different unlearning subsets. Each block (Random, Office-Home Top, CUB Top, DomainNet Top) indicates the subset of classes removed. For each subset, we report $k$-NN Accuracy on three downstream tasks (Office-Home, CUB, and DomainNet-126) to measure transferability. The blue numbers indicate the gap relative to the retrained model. The bold numbers represent optimal values.
Random Office-Home Top CUB Top DomainNet Top
Classes Method Office-Home CKA CUB CKA DomainNet CKA Avg. Office-Home CKA CUB CKA DomainNet CKA
@@
Table 14: Comparison of CKA values between $\theta_{u}$ and $\theta_{r}$ across different unlearning subsets. Each block (Random, Office-Home Top, CUB Top, DomainNet Top) indicates the subset of classes removed. For each subset, we report CKA value on three downstream tasks (Office-Home, CUB, and DomainNet-126) to gauge representational quality. The bold numbers represent optimal values.
Random Office-Home Top CUB Top DomainNet Top
Classes Method AGL AGR H-LR AGL AGR H-LR AGL AGR H-LR AGL AGR H-LR